OpenStudy (a1234):

A rectangular object has a width of 40 meters, height of 15 meters, and length of 2 meters. It floats consistently when 3 meters of its height is below the surface of the water. 1. Find the volume of the displaced water. 2. How much is the buoyant force on the object? 3. What is the object's weight? 4. What is the mass density of the block?

OpenStudy (a1234):

Is the volume of the displaced water equal to 3*40*2, or 240?

OpenStudy (a1234):

And I got 240*1000*9.8 = 2352000 N for the buoyant force

OpenStudy (a1234):

I'm not sure how to get the third and fourth parts.

ganeshie8 (ganeshie8):

Let me ask you a silly question, what is 2352000 N ?

OpenStudy (a1234):

2352000 Newtons, the buoyant force

OpenStudy (a1234):

Density = mass/volume, right? But how do I get the mass with the given information @ganeshie8 ?

ganeshie8 (ganeshie8):

use this weight = mass*g

ganeshie8 (ganeshie8):

I made a silly mistake in part 3. Let's start over.. According to archimedes principle, that must equal the weight of the `COMPLETE object`, yes ?

OpenStudy (a1234):

But that calculation was just for the part that was submerged...what about the rest of the object?

ganeshie8 (ganeshie8):

No, we were wrong about our inerpretation of the archimedes principle earlier. Here is the correct interpretation : The weight of the water displaced equals the weight of the COMPLETE object.

ganeshie8 (ganeshie8):

Let me delete our earlier incorrect discussion, its so embarassing lol

OpenStudy (a1234):

So...what do we do next :)

ganeshie8 (ganeshie8):

So... what's the answer for part 3 ?

OpenStudy (a1234):

2352000 N?

ganeshie8 (ganeshie8):

yes, its same as the part 2 : `Weight of displaced fluid = buoyant force on the object = Weight of the COMPLETE object`

OpenStudy (a1234):

I'm going to use that number in the mass density calculation. mass = weight/g mass = 2352000/9.8 m = 240000 density = mass/volume density = 240000/1200 d = 200

ganeshie8 (ganeshie8):

That looks good to me..

OpenStudy (a1234):

Is it right that I divided by 1200 instead of 240?

ganeshie8 (ganeshie8):

240000 is the mass of the complete object 1200 is the volume of the complete object To get the density, we use the number 1200 yea

OpenStudy (a1234):

Ok...one last question 5. How would the amount of the object under the surface change if it were floating in glycerin? I'm thinking that I should use the density of glycerin instead of water in the calculation for amount of displaced water. Is this right?

ganeshie8 (ganeshie8):

Yes, density of glycerin is greater than the density of water. So I'd guess the height of the submerged part of the object decreases ?

OpenStudy (a1234):

Thank you very much for all the help!

ganeshie8 (ganeshie8):

Np :)